OpenStudy (anonymous):

why is x^2=y^2 not a function but x^3=y^3 is a function?

OpenStudy (freckles):

\[x^2-y^2=0 \text{ \implies } (x-y)(x+y)=0 \\ \text{ this happens when } x=y \text{ or } x=-y \] Therefore x^2=y^2 is not a function... \[x^3-y^3=0 \\\] Try to factor this one

OpenStudy (freckles):

Do you know how to factor a difference of cubes?

OpenStudy (anonymous):

thank u